OpenStudy (anonymous):

Solve the equation 2 cos^2(x) − cos(x) − 1 = 0, 0 ≤ x ≤ 2π.

OpenStudy (mathstudent55):

This is a quadratic equation in cos(x). Substitute y = cos(x), and y^2 = cos^2(x) and solve the quadratic equation for y. Then substitute back cos(x) for y, and solve the trig equation.

OpenStudy (anonymous):

okay, that gives me -1/2, and 1. Now to plug that back into the equation just put those values in for x?

OpenStudy (anonymous):

No. You don't put them in as x. You originally set y = cos(x) so now that you have two y values, solve for what x's does -1/2 = cos(x). Also look at for what x's does 1 = cos(x). Then look at the interval they gave you.

OpenStudy (anonymous):

0, 2pi/3, 4pi/3, and 2pi?

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

awesome, thanks for the help!

OpenStudy (mathstudent55):

\(2 \cos^2x − \cos x − 1 = 0 \) Let \(y = \cos x\) \( 2y^2 - y - 1 = 0\) \( (2y + 1)(y - 1) = 0\) \( 2y + 1 = 0\) or \(y - 1 = 0\) \(2y = -1\) or \(y = 1\) \(y = -\dfrac{1}{2} \) or \(y = 1\) Now we substitute back. \(\color{red}{cos x = -\dfrac{1}{2}} \) or \(\color{green}{\cos x = 1}\) \(\color{red}{x = \dfrac{2\pi}{3}} \) or \(\color{red}{x = \dfrac{4\pi}{3}} \) or \(\color{green}{x = 0} \) or \( \color{green}{x = 2\pi } \) I used red and green to show which solutions come from which equations.